five times the diffrence of a number and one yields the same result as twice the same number increase by three: five times the diffrence of a number and one yields the same result as twice the same number increase by three
Answer by orca(336)

We need to set up an equation.
Let x be the number.
The diffrence of a number and one is x - 1.
Five times the diffrence of a number and one is 5(x - 1)
Twice the same number is 2x.
Twice the same number increase by three is 2x +3
So
5(x - 1) = 2x +3
Solving for x we obtain:
5x - 5 = 2x + 3
3x = 8
x = 8/3
Therefore the number is 8/3.